OpenStudy (anonymous):

Christina works part-time from home as a data entry professional. She enters data in forms and is paid based on the entry accuracy rate. The table below shows her compensation. A 2-column table with 'Accuracy' and 'Compensation per Form' as the column headings. Tier 1 is $15, Tier 2 is $12, and Tier 3 is $0 Based on her regular performance, the equation below can be used to determine C, the total amount Christina earns a month entering data, where x represents the total number of forms she completes that month. C = 15(0.8x) + 12(0.1x) + 0(0.1x)

OpenStudy (anonymous):

What is the minimum number of forms she must complete in a month to earn a total of at least $1,584? I'm thinking C is for $1,584

OpenStudy (anonymous):

thanks for no help . lel 1584=15(0.8x)+12(0.1x)+0(0.1x) 1584=12x+1.2x+x 1584=13.2x 1584/13.2 = 13.2/13.2 x=120

zepdrix (zepdrix):

Hah you're hiding in the bushes! :3 Hmmmm, ya it looks like you did that correctly. Good job! \c:/

zepdrix (zepdrix):

That second line should have been: 1584=12x+1.2x+0x But looks like that's what you meant anyway.

OpenStudy (anonymous):

thank you . cx i was confused if i should have subbed the 1583 for c . & i tried it & it worked . c: can you fan me ? you seem to understand algebra a lot more than i do !

zepdrix (zepdrix):

fan? :O i dont know if ive tried that before